Solve the following equations and check your results. 5t – 3 = 3t – 5 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

5t–3=3t–5
⇒ 5t- 3t = -5 + 3
⇒ 2t =-2
⇒ t =-1
Putting the value of t in RHS and LHS we get,
5× (-1) - 3 = 3× (-1) - 5
⇒-5-3=-3-5
⇒-8=-8
⇒LHS=RHS

Thanks